# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ Abstract
We consider a Moran-type model of cultural evolution, which describes how traits emerge, are transmitted, and get lost in populations. Our analysis focuses on the underlying cultural genealogies; they were first described by Aguilar and Ghirlanda (2015) and are closely related to the ancestral selection graph of population genetics, wherefore we call them ancestral learning graphs. We investigate their dynamical behaviour, that is, we are concerned with evolving genealogies. In particular, we consider the total length of the genealogy of a sample of individuals from a stationary population as a function of the (forward) time at which the sample is taken. This quantity shows a sawtooth-like dynamics with linear increase interrupted by collapses to near-zero at random times. We relate this to the metastable behaviour of the stochastic logistic model, which describes the evolution of the number of ancestors, or equivalently, the number of descendants of a given sample.
We assume that new inventions appear independently in every individual, and all traits of the cultural parent are transmitted to the learner in any given learning event. The set of traits of an individual then agrees with the set of innovations along its genealogy. The properties of the genealogy thus translate into the properties of the trait set of a sample. In particular, the moments of the number of traits are obtained from the moments of the total length of the genealogy.
@@
cultural evolution , ancestral learning graph , ancestral selection graph , stochastic logistic model , metastability , coalescence
@@ 1 Introduction
Stochastic processes in cultural evolution describe how traits emerge, are transmitted, and get lost in populations under certain random laws. This relatively new topic of research bears many connections with population genetics with respect to the models used, the questions asked, and the methods employed. For example, Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018) studied a discrete-time Moran model of cultural evolution under various assumptions on the origination and transmission of traits; they investigated quantities such as the number of traits maintained in a population and the corresponding popularity spectrum, that is, the histogram of trait frequencies in the population, akin to the site-frequency spectrum in population genetics. These and similar results are centered around the traits, focus on the stationary state, and characterise it by means of expectations of various quantities related to trait frequencies.
@@
Aguilar and Ghirlanda (2015) were, to the best of our knowledge, the first to explicitly consider the cultural genealogies in this kind of model. They worked with a continuous-time Moran-type model of cultural evolution, where the genealogy is closely related to the ancestral selection graph (ASG) of population genetics (Neuhauser and Krone, 1997; Krone and Neuhauser, 1997). Aguilar and Ghirlanda (2015) investigated the expected time to the most recent unique ancestor (MRUA), a quantity different from both the most recent common ancestor (MRCA) and the ultimate ancestor (UA) of the ASG. Kobayashi et al. (2018) worked with a discrete-time Wright–Fisher-type model of cultural evolution assuming independent origination and transmission of traits and used genealogical thinking to obtain the expected number and age of distinct cultural traits in a finite sample. Altogether, little is known beyond expectations.
@@
The goal of this paper is threefold. First, we will focus on the underlying cultural genealogies, which are also of independent interest. Second, we aim at more details about the (stationary) distribution of the number of traits; that is, we look at the variance and higher moments. Third, we consider the temporal fluctuations of the number of traits at stationarity. All properties of the trait frequencies are closely tied to the genealogies as well. Exploiting the connection to recent variants of the ASG of population genetics and to the stochastic logistic model of epidemiology and theoretical ecology, we analyse both the mean behaviour and the dynamics. A special role will be played by the concept of evolving genealogies and the complementary descendant process.
@@
The paper is organised as follows. In Section 2, we describe and define our model forward in time and then introduce the genealogical process, which we term the ancestral learning graph (ALG), and elaborate on the connection to ancestral structures in population genetics. In Section 3, we present the results. We first show that the moments of the number of traits in a sample at stationarity can be expressed in terms of the moments of the total tree length of the ALG, that is, the sum of the lengths of all its branches. For the mean and variance, we derive explicit expressions. In simulations, we explore the dynamics of the tree length (and hence the number of traits) as a function of time. We observe an unexpeced sawtooth-like behaviour, which cannot be explained by the moments alone. We understand it by formulating the concept of evolving genealogies (backward in time) and the descendant process (forward in time) and investigating them with the help of the properties of the stochastic logistic process. In particular, the metastable state of the latter and the coalescence of sets of lines in the ALG will play crucial roles. As an aside, we also obtain the popularity spectrum. We discuss our findings in Section 4.
@@ 2 The model
We follow the model of Aguilar and Ghirlanda (2015), who assume a population of fixed size $N$ in continuous time with death-birth and learning events, akin to a Moran model in population genetics. Newborn individuals do not have any cultural parent at the time of birth, but acquire role models one by one through learning events, each of which assigns a (single) new cultural parent to the focal individual; it is intentionally left unspecified which traits are transmitted in a given learning event, and whether and how new traits emerge. We combine this model with a simplistic model that makes the appearance and transmission of traits explicit. This is motivated by the work of Kobayashi et al. (2018), who assume a discrete-time Wright–Fisher process, where every newborn invents a random number of new traits and is assigned a random number $K$ of potential role models; every trait of each of the role models is independently transmitted to the newborn with probability $b$. Here we assume that, within the continuous-time model of Aguilar and Ghirlanda (2015), new traits appear at constant rate $\mu$ in every individual, and we specify the mode of transmission in accordance with Kobayashi et al. (2018) with the choice $b=1$, so that the learning individual acquires all traits of its cultural parent(s). The resulting model may also be considered as a continuous-time version of (special cases of) the models of Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018).
@@ 2.1 The forward process
Consider a population of $N$ individuals. From the Moran model, we borrow the graphical representation shown in Figures 1 and 2. Each individual corresponds to a vertical line segment, with the forward direction of time being top to bottom; the lines have labels in $[N]\mathrel{\mathop{:}}=\{1,2,\ldots,N\}$. The events described below are represented by graphical elements juxtaposed to this picture. Again in line with a common strategy for the Moran model, we describe the model in two steps: first the so-called untyped model, which contains the death-birth and learning events; and second the typed model, which includes the traits, specifically, the innovation events and the transmission of traits on the occasion of learning events.
@@
    Untyped version. Each individual can experience two kinds of events at any instant of time and independently of the remaining population: a death event and a learning event. Death occurs at rate $u>0$ per individual, and when this happens, a newborn will replace it. Learning events occur at rate $s>0$ per individual. In a learning event, the individual randomly samples a cultural parent uniformly from among the $N$ individuals (including itself). So far, the learning events are abstract; we do not yet say what is transmitted along them. In the graphical representation (Fig. 1), death-birth events are symbolised by crosses; they appear at rate $u$ on every line, by way of independent Poisson point processes. Learning events are depicted by arrows between the lines, with the cultural parent at the tail and the learner at the tip; arrows appear at rate $s/N$ per ordered pair of lines, again by way of independent Poisson point processes. We omit arrows where the parent and the learner are identical.
    Figure 1: Graphical representation of a realisation of the untyped model ($N=5$).
@@
    Including the types. The type of an individual is the collection of its cultural traits. Given a realisation of the untyped system, we turn it into a typed one as folllows. We assume that every individual independently invents new traits at rate $\mu>0$. In the graphical representation, these innovations are indicated by circles, which are laid down on every line by way of a Poisson point process at rate $\mu$, see Figure 2. We assume that, first, all innovations are non-recurrent, that is, have never occurred in the population before. Second, in every learning event, the learner acquires all traits carried by the cultural parent. Third, we assume that death events eliminate all traits, so the newborn is devoid of any trait.
    We now assign a set of traits to each line at time $t=0$ in an exchangeable way (that is, according to a law that is invariant under permutation of lines); this set may or may not be empty. If we then propagate the traits forward in time according to the rules just described, we get the types, that is, the collection of traits, for every individual and any time $t>0$.
Figure 2: Graphical representation of a realisation of the typed model.
@@ 2.2 Formal definition of the forward process
$\Phi(t)\mathrel{\mathop{:}}=(K_{1}(t),\ldots,K_{N}(t),\Omega(t))$ (1)
be the state of the typed graphical construction at time $t$, where $K_{\alpha}(t)\subset\mathbb{N}$ denotes the set of traits that individual $\alpha\in[N]$ knows at time $t\in\mathbb{R}_{\geq 0}$, and $\Omega(t)\in\mathbb{N}_{0}$ denotes the number of traits ever invented until $t$, including those already present at time 0; so $\Omega(0)=\lvert\cup_{\alpha\in[N]}K_{\alpha}(0)\rvert$ and $K_{\alpha}(t)\subseteq\{1,2,\dots,\Omega(t)\}$ (where the latter is agreed to be $\varnothing$ if $\Omega(t)=0$), and
@@
note that nothing happens if $\alpha=\beta$. Note also that we have defined the process in a way that is close to how we will later simulate it; in this formulation, it does not become stationary for $t\to\infty$ since, due to the consecutive labelling of the traits, both $(\Omega(t))_{t\geq 0}$ and the $(K_{\alpha}(t))_{t\geq 0}$ are transient. The quantities we will analyse, however, will only rely on the number of traits rather than their labels and do become stationary, as will become obvious in the backward picture.
For a subset of individuals $G\subseteq[N]$, we define their knowledge set as
$K_{G}(t):=\bigcup_{\alpha\in G}K_{\alpha}(t)$ (4)
by slight abuse of notation (so $K_{\{\alpha\}}(t)=K_{\alpha}(t)$). With this, we define the number of traits of a ‘typical’ group of $n$ individuals at time $t$ as
$C_{n}(t)=|K_{[n]}(t)|.$ (5)
By ‘typical’, we here allude to the fact that, by exchangeability, the $K_{G}(t)$ are identically distributed for all $G\subseteq[N]$ with $\lvert G\rvert=n$, so we choose $G=[n]$ as their representative. We will see below that the distribution of $(C_{n}(t))_{t\geq 0}$ will become stationary as $t\to\infty$; we will denote by $C_{n}=C_{n}(\infty)$ a random variable that has this stationary distribution.
@@ 2.3 The ancestral learning graph (ALG)
Figure 3: The ALG for the realisation of the forward model in Figure 2, starting from $t=t_{\text{max}}$. Red and black lines are ancestral and non-ancestral, respectively, to the sample of size $n=2$ taken at $\tau=0$.
@@
    Start the graph by tracing back the lines emerging from the individuals sampled at forward time $t$. Denote backward time by $\tau$, so backward time $\tau$ corresponds to forward time $t-\tau$; in particular, forward times 00 and $t$ correspond to backward times $\tau=t$ and $\tau=0$, respectively. Proceed as follows in an iterative way in the backward direction of time until the initial time is reached, or until there are no lines left.
@@
      If a line currently in the graph is hit by the tip of a learning arrow, we trace back both ancestors, namely that of the cultural parent (at the tail of the arrow, also known as the incoming branch) and that of the learning individual (at the tip; also known as the continuing branch). That is, we add the incoming line to the graph; this results in a branching event. Note that the graph remains unchanged if the incoming branch is already contained in it.
@@
      If a line currently in the graph is hit by a cross, we do not follow it back any further, that is, we prune it.
    The resulting untyped ALG consists of all cultural ancestors of the sampled individuals over time, including the individuals in the sample themselves until they die (that is, individuals are counted as their own cultural ancestors); the untyped ALG corresponds to the untyped version of the forward model.
@@
    The untyped ALG may be turned into a typed one by first assigning, to each line in the graph that is still alive at backward time $t$, the initial trait set from the Moran model; if there are no lines left at backward time $t$, this step is void. Then, the traits are propagated forward (that is, downward) in the genealogy according to the same rules as before. That is, when a new trait was acquired in the Moran model on a line belonging to the genealogy, it is added to the trait set of the individual; in each branching event, the trait set of the parent is added to that of the learner; if a line encounters a cross, it loses all traits. This way, a trait set is associated with every line element of the graph.
Note that the untyped ALG obtained in Step 1 is the cultural genealogy of Aguilar and Ghirlanda (2015), with the only difference that we allow for individuals to also choose themselves as cultural parents; but this is an event without any effect, so it only introduces a tiny change of time scale. The typing that happens in Step 2 describes how traits appear and are transmitted along the given genealogy; this step was intentionally left unspecified by Aguilar and Ghirlanda (2015).

### v2

# Evolving genealogies in cultural evolution, the descendant process, and the number of cultural traits

@@ Abstract
We consider a Moran-type model of cultural evolution, which describes how traits emerge, are transmitted, and get lost in populations. Our analysis focuses on the underlying cultural genealogies; they were first described by Aguilar and Ghirlanda (2015) and are closely related to the ancestral selection graph of population genetics, wherefore we call them ancestral learning graphs. We investigate their dynamical behaviour, that is, we are concerned with evolving genealogies. In particular, we consider the total length of the genealogy of the entire population as a function of the (forward) time where we start looking back. This quantity shows a sawtooth-like dynamics with linear increase interrupted by collapses to near-zero at random times. We relate this to the metastable behaviour of the stochastic logistic model, which describes the evolution of the number of ancestors as well as the number of descendants of a given sample.
We superpose types to the model by assuming that new inventions appear independently in every individual, and all traits of the cultural parent are transmitted to the learner in any given learning event. The set of traits of an individual then agrees with the set of innovations along its genealogy. The properties of the genealogy thus translate into the properties of the trait set of a sample. In particular, the moments of the number of traits are obtained from the moments of the total length of the genealogy.
@@
cultural evolution , ancestral learning graph , ancestral selection graph , stochastic logistic model , metastability , merging of genealogies
@@ 1 Introduction
Cultural traits originate by innovation and are transmitted by learning, as opposed to genetic traits, which originate by mutation and are passed on via inheritance. Cultural evolution is about how cultural traits emerge, are transmitted, and lost in populations. The quantitative description and mathematical modelling of cultural evolution has a venerable history, as testified, for example, by the classical monographs of Cavalli-Sforza and Feldman (1981) and Boyd and Richerson (1985). Stochastic models of cultural evolution are, however, relatively recent. Here, models of population genetics are often used as blueprints, also with respect to the questions asked and the methods employed. For example, Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018) studied a discrete-time Moran-type model of cultural evolution under various assumptions on the origination and transmission of traits; they investigated quantities such as the number of traits maintained in a population and the corresponding popularity spectrum, that is, the histogram of trait frequencies in the population, akin to the site-frequency spectrum in population genetics. These and similar results are centered around the traits, focus on the stationary state, and characterise it by means of expectations of various quantities related to trait frequencies.
@@
Aguilar and Ghirlanda (2015) were, to the best of our knowledge, the first to explicitly consider the cultural genealogies in this kind of model. They worked with a continuous-time Moran-type model of cultural evolution, where the genealogy is closely related to the ancestral selection graph (ASG) of population genetics (Neuhauser and Krone (1997); Krone and Neuhauser (1997); see also Mano (2009)). Aguilar and Ghirlanda (2015) investigated the expected time to the most recent unique ancestor (MRUA), a quantity different from both the most recent common ancestor (MRCA) and the ultimate ancestor (UA) of the ASG. Kobayashi et al. (2018) worked with a discrete-time Wright–Fisher-type model of cultural evolution assuming independent origination and transmission of traits and used genealogical thinking to obtain the expected number and age of distinct cultural traits in a finite sample. Altogether, little is known beyond expectations.
@@
The goal of this paper is twofold. We will first focus on the cultural genealogies, which are of independent interest. In particular, we consider the total length of the genealogies (that is, the sum of the lengths of all its branches) as a function of the time where we start looking back, that is, we consider evolving genealogies. Exploiting the connection to recent variants of the ASG of population genetics and to the stochastic logistic model of epidemiology and theoretical ecology, we analyse both the moments and the temporal fluctuations of the length of the genealogies at stationarity. Second, we superpose the traits on the model via admittedly simplistic assumptions, namely 1) traits originate independently, 2) all traits of a cultural parent are jointly transmitted to the learner, and 3) traits are lost at death only. This way, the distribution of the number of traits is closely tied to the genealogies, but contains additional randomness.
@@
The paper is organised as follows. In Section 2, we describe and define our model forward in time and then introduce the genealogical process, which we term the ancestral learning graph (ALG); we also elaborate on the connections to population genetics and epidemiology. In Section 3, we present the results. We first consider the moments of both the length of the genealogy and the number of traits at stationarity, and then turn to their distributions. From there, we move on to the dynamics of the ALG, first when started from a fixed time and then as a function of the forward time where we start looking back. In simulations, we explore the dynamics of the length of the resulting evolving genealogy (and hence the number of traits) and observe a surprising sawtooth-like behaviour. We understand it by investigating the evolving genealogies (backward in time) and the corresponding descendant process (forward in time) with the help of the properties of the stochastic logistic process. In particular, the metastable state of the latter and the merging of ALGs will play crucial roles. As an aside, we also obtain the popularity spectrum. We discuss our findings in Section 4.
@@ 2 The model
We follow the model of Aguilar and Ghirlanda (2015), who assume a population of fixed size $N$ in continuous time with death-birth and learning events. We primarily think of individuals as humans, but they could be reinterpreted as groups of humans, see the discussion. The model bears elements of both the Moran model of population genetics and the SIS (susceptible-infected-susceptible) model of epidemiology; we will discuss these connections later. We assume that newborn individuals do not have any cultural parent at the time of birth, but acquire role models one by one through learning events, each of which assigns a (single) new cultural parent to the focal individual. Aguilar and Ghirlanda (2015) restrict themselves to this so-called untyped model, where it is intentionally left unspecified which traits are transmitted in a given learning event, and whether and how new traits emerge. We combine this with a simplistic model that makes the appearance and transmission of traits explicit. This is motivated by the work of Kobayashi et al. (2018), who assume a discrete-time Wright–Fisher process, where every newborn invents a random number of new traits and is assigned a random number $K$ of potential role models; every trait of each of the role models is independently transmitted to the newborn with probability $b$. Here we assume that, within the continuous-time model of Aguilar and Ghirlanda (2015), new traits appear at constant rate $\mu$ in every individual, and we specify the mode of transmission in accordance with Kobayashi et al. (2018) with the choice $b=1$, so the learning individual acquires all traits of its cultural parent; this assumption is made for mathematical convenience. The resulting model may also be considered as a continuous-time version of (special cases of) the models of Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018). In the discussion, we will — as an outlook — waive the admittedly unrealistic assumption of all-or-none transmission and explore what happens if every trait of the role model is independently transmitted to the learner with probability $b<1$.
@@ 2.1 The forward process
Consider a population of $N$ individuals. We will work with the graphical representation as introduced by Harris (1978) (see Liggett (2011) for a review). This is now an indispensable tool in stochastic models in biology and physics and, for our case, shown in Figures 1 and 2. Each individual corresponds to a horizontal line segment, with the forward direction of time being left to right; the lines have labels in $N]\mathrel{\mathop{:}}=\{1,2,\ldots,N\}$ and extend through all of $\mathbb{R}$. The events described below are represented by graphical elements superposed to this picture. Again in line with a common strategy for the Moran model, we describe the model in two steps: first the untyped model, which contains the death-birth and learning events; and second the typed model, which includes the traits, specifically, the innovation events and the transmission of traits on the occasion of learning events.
@@
    Untyped version. Every individual can experience two kinds of events at any instant of time and independently of the remaining population: a death event and a learning event. Death occurs at rate $u>0$ per individual, and when this happens, a newborn will replace it. Learning events occur at rate $s>0$ per individual. In a learning event, the individual randomly samples a cultural parent uniformly from among the $N$ individuals (including itself). So far, the learning events are abstract; we do not yet say what is transmitted along them. In the graphical representation (Fig. 1), death-birth events are symbolised by crosses; they appear at rate $u$ on every line, by way of independent Poisson point processes. Learning events are depicted by arrows between the lines, with the cultural parent at the tail and the learner at the tip; arrows appear at rate $s/N$ per ordered pair of lines, again by way of independent Poisson point processes on $\mathbb{R}$. We omit arrows where the parent and the learner are identical.
    Figure 1: Graphical representation of a realisation of the untyped model ($N=5$). The lines extend through all of $\mathbb{R}$ and are decorated with crosses (death-birth events) and learning arrows. In this way, the time evolution is defined for any time interval such as $0,t_{\text{max}}]$.
@@
    Including the types. The type of an individual is the collection of its cultural traits. Given a realisation of the untyped system, we turn it into a typed one as folllows. We assume that every individual independently invents new traits at rate $\mu>0$. In the graphical representation, these innovations are indicated by circles, which are laid down on every line by way of a Poisson point process on $\mathbb{R}$ at rate $\mu$, see Figure 2. We assume that, first, all innovations are non-recurrent, that is, have never occurred in the population before. Second, in every learning event, the learner acquires all traits carried by the cultural parent; this happens by merging the knowledge sets of the learner and the parent. Third, we assume that death events eliminate all traits, so the newborn is devoid of any trait.
    We now assign a set of traits to every line at time $t=0$ in an exchangeable way (that is, according to a law that is invariant under permutation of lines); this set may or may not be empty. From $t=0$ onwards, we then number the innovations consecutively in the order of their appearance. If we finally propagate the traits forward in time according to the rules just described, we get the types, that is, the collection of traits, for every individual and any time $t>0$. Let now $K_{\alpha}(t)\subset\mathbb{N}$ be the set of traits that individual $\alpha\in[N]$ knows at time $t\in\mathbb{R}_{\geq 0}$; that is, the individual’s knowledge set. For a subset of individuals $G\subseteq[N]$, we define
    $K_{G}(t)\mathrel{\mathop{:}}=\bigcup_{\alpha\in G}K_{\alpha}(t)$ (1)
    by slight abuse of notation (so $K_{\{\alpha\}}(t)=K_{\alpha}(t)$). With this, we denote the number of traits of a ‘typical’ group of $n$ individuals at time $t$, that is, the cardinality of the group’s knowledge set, as
    $C_{n}(t)\mathrel{\mathop{:}}=|K_{[n]}(t)|.$ (2)
    By ‘typical’, we here allude to the fact that, by exchangeability, the $K_{G}(t)$ are identically distributed for all $G\subseteq[N]$ with $\lvert G\rvert=n$, so we choose $G=[n]$ as their representative. Clearly, then, $C_{n+1}(t)\geq C_{n}(t)$ for all $t$. We will see below that the distribution of $(C_{n}(t))_{t\geq 0}$ will become stationary as $t\to\infty$; we will denote by $C_{n}=C_{n}(\infty)$ a random variable that has this stationary distribution.
Figure 2: A typed version of Figure 1 in the interval $0,t_{\text{max}}]$.
@@ 2.2 Formal definition of the forward process
$\Phi(t)\mathrel{\mathop{:}}=\big(K_{1}(t),\ldots,K_{N}(t),\Omega(t)\big)$ (3)
be the state of the typed graphical construction at time $t$, where $\Omega(t)\in\mathbb{N}_{0}$ denotes the number of traits ever invented until $t$, including those assigned at time 0; so $\Omega(0)=\lvert\cup_{\alpha\in[N]}K_{\alpha}(0)\rvert$ and $K_{\alpha}(t)\subseteq\{1,2,\dots,\Omega(t)\}$ (where the latter is agreed to be $\varnothing$ if $\Omega(t)=0$), and
@@
note that nothing happens if $\alpha=\beta$. Note also that we have defined the process in a way that is close to how we will later simulate it; in this formulation, it does not become stationary for $t\to\infty$ since, due to the consecutive labelling of the traits, both $(\Omega(t))_{t\geq 0}$ and the $(K_{\alpha}(t))_{t\geq 0}$ are transient. The quantities we will analyse, however, will only rely on the number of traits rather than their labels and converge to stationarity as $t\rightarrow\infty$, as will become obvious in the backward picture.
@@ 2.3 The ancestral learning graph (ALG)
Figure 3: An ALG embedded in the realisation of the forward model in Figure 2, starting from $t=t_{\text{max}}$. Red and black lines are ancestral and non-ancestral, respectively, to the sample of size $n=2$ taken at $\tau=0$.
@@
    Start the graph by tracing back the lines emerging from the individuals sampled at forward time $t$. Denote backward time by $\tau$, so backward time $\tau$ corresponds to forward time $t-\tau$; in particular, forward times 0 and $t$ correspond to backward times $\tau=t$ and $\tau=0$, respectively. Proceed as follows in an iterative way in the backward direction of time until the initial time is reached, or until all ancestral lineages have been hit by a cross.
@@
      If a line currently in the graph is hit by the tip of a learning arrow, we trace back both ancestors, namely that of the cultural parent (at the tail of the arrow; in line with ASG terminology, we call it the incoming branch) and that of the learning individual (at the tip; again, as in the ASG, we call it the continuing branch). That is, we add the incoming line to the graph; this results in a branching event. Note that the graph remains unchanged if it already contains the incoming branch.
@@
      If a line currently in the graph is hit by a cross, we do not follow it any further, that is, we prune it.
    The resulting untyped ALG consists of all cultural ancestors of the sampled individuals over time, including the individuals in the sample themselves until their births (that is, at branching events, individuals are counted as their own cultural ancestors); the untyped ALG corresponds to the untyped version of the forward model.
@@
    The untyped ALG may be turned into a typed one by first assigning, to each line in the graph that is still alive at backward time $t$, the initial trait set from the forward model; if there are no lines left at backward time $t$, this step is void. Then, the traits are propagated forward (that is, to the right) in the genealogy according to the same rules as before. That is, when a new trait was acquired in the forward model on a line belonging to the genealogy, it is added to the trait set of the individual; in each branching event, the trait set of the parent is united with that of the learner; and if a line encounters a cross, it loses all traits. This way, a trait set is associated with every line element of the graph.
Note that the untyped ALG obtained in Step 1 is the cultural genealogy of Aguilar and Ghirlanda (2015), with the only difference that we allow for individuals to also choose themselves as cultural parents; but this is an event without any effect, so it only changes the time scale by a factor of (N1)/N(N-1)/N. The typing in Step 2 describes how traits appear and are transmitted along the given genealogy; this step was intentionally left unspecified by Aguilar and Ghirlanda (2015). Note also that we can obtain a stationary situation for any $t\geq 0$ by running the untyped ALG until all (ancestral) lines have been hit by a cross, even if this requires $\tau>t$; this is where the graphical construction on all of $\mathbb{R}$ comes in handy. Turning the untyped ALG into a typed one then does not require the assignment of initial types — the types are solely determined by origination and transmission of traits along the graph.